Circular No. 6459

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1996 N2 (ELST-PIZARRO) P. Pravec, Ondřejov Observatory, writes: “Measurements of my CCD

images of P/1996 N2 taken with the 0.65-m telescope on Aug. 21.973– 21.977 UT reveal the presence of a stellar nucleus and a straight, narrow tail 3′.5 long extending from the nucleus along a line in p.a. 251^{o}.4 ± 0^{o}.4 with dispersion of a few arcsec. This is only 1^{o} less than the p.a. of the line of variation of T for the orbit published on IAUC 6457, but it is nearly opposite the antisolar direction. If the tail is composed of material spread along the orbit, the material is distributed within an interval of ΔT from 0.00 to +0.15 day with respect to the object’s position. However, the projection of a relatively normal dust tail could cause a similar phenomenon, if the tail deviates from the prolonged radius vector, because the earth is very close to the object’s orbital plane.”

Z. Sekanina, Jet Propulsion Laboratory, reports: “The comet’s narrow, straight, structureless tail is likely to be a signature of a past dust-emission episode. Even though the very low inclination strongly degrades the accuracy with which the event’s timing can be determined, the 1σ error of the tail’s position angle reported by Pravec (above) suggests that the presumed outburst probably occurred between late May and early July 1996, or some 40 to 80 days after perihelion, in which case the maximum effect of solar radiation pressure on dust in the tail was between 0.05 and ∼0.4 the solar attraction (implying the presence of micron- or submicron-sized grains). Information from the Elst tail-length estimate in mid-July and from Offutt’s observation on Aug. 21 (IAUC 6456) leads to similar constraints. In an effort to improve the time determination of the presumed outburst, it is desirable to measure the tail’s position angle with maximum possible accuracy. It is also recommended that a search be initiated for possible prediscovery observations from the critical interval to ascertain the object’s physical appearance at the time.”

GRS 1915+105 S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on

behalf of the GRANAT/WATCH team, report x-ray observations on June 7, simultaneous with those reported on IAUC 6455, when GRS 1915+105 was in a high flaring state, with a flux of ∼ 0.4 Crab in the range 8–20 keV.

1996 August 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green